# Classically-embedded split Cayley hexagons rule three-qubit contextuality with three-element contexts

### v1

@@ 2 Contextual configurations in symplectic polar spaces, degree of contextuality
Figure 4: A Mermin-Peres magic square, aka $\mathcal{Q}^{+}(3,2)$, and the doily, aka $\mathcal{W}(3,2)$, are two ‘extremal’ two-qubit observable-based contextual configurations. The minimal number of constraints that cannot be satisfied by an NCHV model, i. e. the degree of contextuality, is one for the grid and three for the doily. For these ‘peculiar’ configurations, with a single exception, the unsatisfiable constraints can be identified solely with the corresponding negative contexts.
@@ 3 The two types of symplectic embeddings of the smallest split Cayley hexagon
Although the two-qubit symplectic space $\mathcal{W}(3,2)$ already contains the fundamental building blocks furnishing observable-based proofs of quantum contextuality, namely the above-discussed grids as well as two-spreads [muller2023new], it is the three-qubit space $\mathcal{W}(5,2)$ where the power of our formalism acquires a completely new dimension. This is mainly because, as already briefly described in Section 2, this space, in addition to quadrics, features also another distinguished subgeometry – the split Cayley hexagon of order two, $\mathcal{H}$, and in its two non-isomorphic embeddings at that. Recently three of us [holweck2022three] discovered that the two unequivalent embeddings of this hexagon into $\mathcal{W}(5,2)$ behave differently when it comes to their line-complements; in particular, it was demonstrated that the complement of any skew-embedded copy of $\mathcal{H}$ is contextual, which is, strangely, not the case for any classically-embedded one. In what follows we will demonstrate that classically-embedded copies of $\mathcal{H}$ also enter the game, but in a different and rather unexpected way. To this end in view, it is necessary to address first in more detail the principal difference between the two embeddings.
@@ 3.1 Classical versus skew embeddings of the hexagon
To see the fundamental difference between the two embeddings, let us call a point of the split Cayley hexagon of order two located in $\mathcal{W}(5,2)$ planar if all the three lines passing through it lie in a plane of $\mathcal{W}(5,2)$. A classically embedded hexagon enjoys the property that each of its points is planar. In a skew embedded hexagon, however, there are only 15 points that are planar; they lie on six lines forming three concurrent pairs, the three points of concurrence lying on the axis of the hexagon. For each of the remaining 48 points only two lines passing through it lie in a plane of $\mathcal{W}(5,2)$.
@@
here, the first seven planes are perp-planes, with the underlined first elements being the nuclei/centers of the corresponding perp-sets. It is also worth adding that through any two Heawood planes sitting in the same Heawood graph there pass altogether eight spreads of this type.
Figure 5: An illustration of a spread of planes of the first kind. The individual planes (as point sets) are distinguished from each other by different colors, the two planes of Heawood type being represented by black and white (empty) bullets; the nuclei/centers of the perp-sets are encircled. This drawing of the split Cayley hexagon of order two, in a form showing its automorphism of order seven, is a slight modification of those given in [schr, psm]. Labeling by the three-qubit observables is taken from [lsv]; see also [sppl, psh].
@@
In a skew embedded hexagon, however, neither of the above-described two patterns can be found; this is mainly due to an insufficient number of planar points, but also due to the way how these points are arranged with respect to each other.
3.2 Skew-embedded hexagon, linear doilies and
contextuality
@@ 3.2 Skew-embedded hexagon, linear doilies and
The fact that a copy of the split Cayley hexagon of order two embedded skewly into $\mathcal{W}(5,2)$ features non-planar points has a number of interesting consequences. We will briefly address one of them. Given a skew embedded hexagon, like the one depicted in Figure 7, let us pick up in it a line that consists solely of non-planar points, for example the line $IIZ-XIZ-XII$ (drawn black in Figure 7). Through each point of this line there pass one line that does not belong to the plane of $\mathcal{W}(5,2)$ defined by the other two concurrent lines at this point; these are the lines $IIZ-YZZ-YZI$ (via point $IIZ$ – red), $XIZ-YYY-ZYX$ (via $XIZ$ – yellow) and $XII-XXY-IXY$ (via $XII$ – blue). Clearly, these three lines are pairwise disjoint as otherwise our generalized hexagon would contain triangles, which is impossible as its smallest ordinary polygons are hexagons. These three lines define a unique linear doily, namely the one depicted at the bottom of Figure 7. However, this doily shares with our hexagon two more lines (shown in gray) that are disjoint from each other and also from any of the three lines, i. e. they form with the three lines a spread of lines of the doily. It is obvious that such a set of six lines is the maximum set of lines a doily and a skew-embedded hexagon can share; indeed, assuming that our doily shares an additional line with the hexagon that is skew to the black line (i. e. not incident with it) would mean that the latter would contain quadrangles, which contradicts its definition.
@@
Figure 7: An illustration of the procedure that shows that to each ‘non-planar’ line (black) of a skew embedded split Cayley hexagon of order two (top) one can associate a unique linear doily (bottom) that shares with it the maximum possible number of lines. The horizontal highlighted line of the hexagon is its axis; the meaning of the remaining colored lines is explained in the text.
@@
The above-described relation can also help us understand why a classically embedded hexagon and a doily share just three lines belonging to a grid of the doily. For if we disregard in Figure 7 the three colored lines that occur only in skew embedded hexagons, the remaining three lines (i. e. the black line and the two gray lines) indeed belong to a particular grid of the doily in question!
@@
By a computer search we have found out that there are only three more patterns of lines a linear doily can share with a skew embedded hexagon. One of them is an already mentioned set of three mutually disjoint lines that belong to some grid of the doily. There exists another three-line pattern, namely that comprising two disjoint lines having a common transversal; slightly rephrased, this pattern entails any three lines forming sides of a quadrangle in the doily. The remaining type features two intersecting lines. These can readily be illustrated employing the copy of skew embedded hexagon depicted in Figure 8. To this end, let us consider three particular linear three-qubit doilies, namely the doilies whose all 15 observables feature $I$ on the first qubit (the ‘left’ doily), on the second qubit (the ‘middle’ doily) and on the third qubit (the ‘right’ doily). From Figure 8 one can easily discern that the ‘left’ doily (red) shares with the hexagon the following three lines $IXX-IZZ-IYY$, $IYY-IZX-IXZ$ and $IXZ-IIZ-IXI$, the middle one being indeed incident with either of the remaining two that are disjoint. On the other hand, the ‘right’ doily (blue) has two concurrent lines in common with the hexagon: $XZI-IZI-XII$ and $IZI-YZI-YII$. For the sake of completeness, we also mention that the ‘middle’ doily (not shown) shares with the hexagon the maximal pattern, composed of the five lines $ZIY-IIY-ZII$, $YII-IIX-YIX$, $ZIX-XIY-YIZ$, $XIX-YIY-ZIZ$ and $XIZ-IIZ-XII$ forming a spread, and the line $ZII-IIX-ZIX$.
@@
Figure 8: A different copy of skew embedded hexagon (whose axis is the line $YYZ-IXY-YZX$) employed to illustrate additional doily-hexagon intersection patterns. The points and (hexagon-shared) lines of the two doilies are highlighted in different colors, the three points that belong to both the doilies being colored gray.
@@
Apart from classification’s results of their own importance, there is another crucial aspect stemming from the above-discussed intersection properties. Let us consider a doily having with a skew embedded hexagon two concurrent lines in common. It is easy to see that out of the ten grids in this doily, there are four of them that do not contain any of the two shared lines. That is, these four grids are fully contained in the complement of the hexagon. A similar situation also occurs in the above-described three-line pattern; in this case the corresponding doily has just two grids devoid any of the three lines and, so, lying fully in the hexagon’s complement. As grids are the simplest, and so fundamental, quantum contextual structures in multi-qubit symplectic polar spaces, their occurrence in the complements of skew embedded hexagons lends itself as one of the most natural justifications why these complements are contextual.
@@ 4 The contextuality of $\mathcal{W}(5,2)$ and the classical embedding of $\mathcal{H}$
Recently, the authors of [muller2023new] made the key (computer-based) discovery that the degree of contextuality of the configuration comprising all 315 line-contexts of $\mathcal{W}(5,2)$ is equal to $63$, and not $90$ (= the number of negative lines) as previously proposed by [Cab10]. In addition, and more importantly, they found out that the 63 unsatisfiable contexts of this configuration are in bijection with 63 lines of a copy of $\mathcal{H}$ that is embedded classically into $\mathcal{W}(5,2)$. These facts came as a big surprise to us and prompted us to have a closer look at what is going on here. We will provide first some algebraic-geometrical arguments for the occurrence of the number $63$. Then, employing a line-layered decomposition of the hexagon, we will demonstrate on several examples how the contextuality properties of a three-qubit configuration having three-element contexts can simply be read off from its generic intersection with a classically-embedded copy of $\mathcal{H}$.
4.1 The degree of contextuality of $\mathcal{W}(5,2)$ is sixty-three
@@ 4.2 Layering of the split Cayley hexagon of order two
Figure 9: An illustration of the line-distribution within the hexagon.
To see finer traits of the structure of the hexagon, one can redraw the three main ‘layers’ or ‘domains’ of the hexagon, namely red (top left), blue (top right) and yellow (bottom) in the most symmetric way. Note that both the red and blue domains can each be viewed as a pair of circumscribed octagons, either of them being isomorphic to the same $(24_{2},16_{3})$-configuration. In the illustration of the yellow domain each observable is represented by two different (opposite) points and the affine part of each yellow line has four distinct images forming a quadrangle. Our option for such a rendering of the yellow layer is simple: if one stacks all the three figures above each other then the corresponding three observables at a given position define a particular gray line of the hexagon; for example, the three topmost observables form the gray line $XXI-ZYX-YZX$.
@@ 4.3 A simple recipe how to get from a skew-embedded hexagon a classically-embedded one, and vice versa
The above-described layered structure of the hexagon turns out to be very relevant to better understand the relation between the two types of embeddings of the split Cayley hexagon of order two. In fact, there exists a rather simple recipe that ‘transforms’ a skew hexagon into a classical one. Let us start with the skew hexagon shown in Figure 9. Keep the three observables of the axis intact. On each yellow line, swap the two remaining observables. On each red (blue) line leave each observable intact and on each blue (red) line replace its observable by that which is the product of a swapped yellow observable and a fixed red (blue) observable on the corresponding gray line. What we get is a copy of a classically embedded hexagon, depicted in Figure 10, that shares 39 lines with the original skew hexagon: the black line, the 6 yellow lines, the 16 blue lines and the 16 red lines.
@@
This construction only works if the black (reference) line is the axis of the skew hexagon. Obviously, we can reverse the process and start with a classically-embedded hexagon to get a skew one. In this case any of its 63 lines can be taken as the reference black line, so we get 63 different skew copies from a given classical one. As there are 120 different classical copies of the hexagon in $\mathcal{W}(5,2)$, the above property also implies that there are as many as $120\times 63=7560$ skew embedded hexagons living in $\mathcal{W}(5,2)$ – confirming the result of [holweck2022three] based on an exhaustive computer search.
4.4 Classically-embedded hexagons encode three-qubit
contextuality

### v3

@@ 2 Contextual configurations in symplectic polar spaces, degree of contextuality
Figure 4: A Mermin-Peres magic square, aka $\mathcal{Q}^{+}(3,2)$, and the doily, aka $\mathcal{W}(3,2)$, are two ‘extremal’ two-qubit observable-based contextual configurations. The minimal number of constraints that cannot be satisfied by an NCHV model, i. e. the degree of contextuality, is one for the grid and three for the doily. For these ‘peculiar’ configurations, with a single exception, the unsatisfiable constraints can be identified solely with the corresponding negative contexts (represented here by the doubled lines).
For example, if one considers the Magic Mermin pentagram in Figure 3, its degree of contextuality is $1$. In the two-qubit case, the Mermin-Peres magic square and the configuration comprising all 15 three-element contexts of two-qubit Pauli matrices, aka $\mathcal{W}(3,2)$, are both contextual with their degree of contextuality being, respectively, $1$ and $3$. Figure 4 illustrates how these configurations can be parametrized by two-qubit observables. The lines/contexts with a negative sign are doubled. It is clear that assigning $+1$ to each node provides a classical model that satisfies all conditions except those imposed by the negative lines. The fact that the degree of contextuality is $1$ and $3$ indicates that one cannot do better with any NCHV model. Note that the grid is a subgeometry of $\mathcal{W}(3,2)$, being isomorphic to $\mathcal{Q}^{+}(3,2)$; it can be shown that there are in fact ten copies of the grid, i. e. the triangle-free configuration with nine points and six lines, with three points on a line and two lines through a point, lying in $\mathcal{W}(3,2)$. These two types of contextual two-qubit configurations have two remarkable properties that, as we will see, are absent in the three-qubit (and very likely in any higher rank) case. The first one is the fact that the unsatisfiable contexts do not cover all the observables of the configuration. The other notable fact is that, save for a single copy of the grid that features three negative contexts, the corresponding unsatisfiable contexts can be identified in each configuration with its negative contexts.
The above-discussed two two-qubit observable-based proofs of the KS Theorem can be considered, together with the Mermin pentagram, as geometrical building blocks of contextuality. Indeed, as recently proved by two of us [28], once a configuration has been found to be contextual for a particular labeling by Pauli observables, then one can deduce that the same geometric configuration will be contextual and will have the same degree of contextuality no matter what admissible multi-qubit Pauli parametrization is employed. In particular, the fact that the grids and $\mathcal{W}(3,2)$ are contextual implies that $\mathcal{W}(2n-1,2)$ is contextual for all $n\geq 2$ because the symplectic polar space of a given dimension always contains copies of symplectic polar spaces of a smaller dimension.
@@ 3 The two types of symplectic embeddings of the smallest split Cayley hexagon
Although the two-qubit symplectic space $\mathcal{W}(3,2)$ already contains the fundamental building blocks furnishing observable-based proofs of quantum contextuality, namely the above-discussed grids as well as two-spreads [29], it is the three-qubit space $\mathcal{W}(5,2)$ where the power of our formalism acquires a completely new dimension. This is mainly because, as already briefly described in Section 2, this space, in addition to quadrics, features also another distinguished subgeometry – the split Cayley hexagon of order two, $\mathcal{H}$, and in its two non-isomorphic embeddings at that. Recently three of us [18] discovered that the two inequivalent embeddings of this hexagon into $\mathcal{W}(5,2)$ behave differently when it comes to their line-complements; in particular, it was demonstrated that the complement of any skew-embedded copy of $\mathcal{H}$ is contextual, which is, strangely, not the case for any classically-embedded one. In what follows we will demonstrate that classically-embedded copies of $\mathcal{H}$ also enter the game, but in a different and rather unexpected way. To this end in view, it is necessary to address first in more detail the principal difference between the two embeddings.
@@ 3.1 Classical versus skew embeddings of the hexagon
To see the fundamental difference between the two embeddings, let us call a point of the split Cayley hexagon of order two located in $\mathcal{W}(5,2)$ planar if all the three lines passing through it lie in a plane of $\mathcal{W}(5,2)$ i. e. the seven observables lying on these three lines mutually commute. A classically-embedded hexagon enjoys the property that each of its points is planar. In a skew-embedded hexagon, however, there are only 15 points that are planar; they lie on six lines forming three concurrent pairs, the three points of concurrence lying on the axis of the hexagon. For each of the remaining 48 points only two lines passing through it lie in a plane of $\mathcal{W}(5,2)$.
To illustrate this difference in more detail, let us consider a copy of the split Cayley hexagon of order two embedded classically in $\mathcal{W}(5,2)$. The 135 planes of the latter space then split into two disjoint, unequally-sized families having 63 and 72 elements. Every plane of the first family originates, as a point set, from the perp-set of a point of the hexagon, i. e. the set of points collinear with a given point (called the center/nucleus), the point itself inclusive (and henceforth referred to as a perp-plane). The planes of the second family form in the hexagon 36 pairs, each pair – together with corresponding parts of lines of the hexagon – representing a copy of the Heawood graph [16] (a Heawood plane); an illustration of such a pair of Heawood planes is given in Figure 5, which employs the most common, seven-fold-symmetric rendering of the Heawood graph.
Figure 5: A copy of the Heawood graph accommodating, as sets of observables, a pair of three-qubit Fano planes, one represented by seven black bullets and the other by seven big circles; also shown are the remaining 21 points (gray) on the lines represented by the 21 edges of the graph.
@@
here, the first seven planes are perp-planes, with the underlined first elements being the nuclei/centers of the corresponding perp-sets.
Figure 6: An illustration of a spread of planes of the first kind. The individual planes (as point sets) are distinguished from each other by different colors, the two planes of Heawood type being represented by black and white (empty) bullets; the nuclei/centers of the perp-sets are encircled. This drawing of the split Cayley hexagon of order two, in a form showing its automorphism of order seven, is a slight modification of those given in [39, 33]; here, a point of the hexagon is represented by a small circle and a line – each accommodating a triple of pairwise commuting observables whose product is proportional to $I^{\otimes 3}$ – either by a straight segment (seven of them) or by an arc of variable size and curvature. Labeling by the three-qubit observables is taken from [25]; see also [38, 32].
@@
In a skew-embedded hexagon, however, neither of the above-described two patterns can be found; this is mainly due to an insufficient number of planar points, but also due to the way how these points are arranged with respect to each other.
3.2 Skew-embedded hexagon, linear doilies and contextuality
@@
The fact that a copy of the split Cayley hexagon of order two embedded skewly into $\mathcal{W}(5,2)$ features non-planar points has a number of interesting consequences. We will briefly address one of them. Given a skew-embedded hexagon, like the one depicted in Figure 8, let us pick up in it a line that consists solely of non-planar points, for example the line $IIZ-XIZ-XII$ (drawn black in Figure 8). Through each point of this line there pass one line that does not belong to the plane of $\mathcal{W}(5,2)$ defined by the other two concurrent lines at this point; these are the lines $IIZ-YZZ-YZI$ (via point $IIZ$ – red), $XIZ-YYY-ZYX$ (via $XIZ$ – yellow) and $XII-XXY-IXY$ (via $XII$ – blue). Clearly, these three lines are pairwise disjoint as otherwise our generalized hexagon would contain triangles, which is impossible as its smallest ordinary polygons are hexagons. These three lines define a unique linear doily (i. e. a doily located in a certain PG$(3,2)$ of the ambient PG$(5,2)$), namely the one depicted at the bottom of Figure 8. However, this doily shares with our hexagon two more lines (shown in gray) that are disjoint from each other and also from any of the three lines. It is obvious that such a set of six lines is the maximum set of lines a doily and a skew-embedded hexagon can share; indeed, assuming that our doily shares an additional line with the hexagon that is skew to the black line (i. e. not incident with it) would mean that the latter would contain quadrangles, which contradicts its definition.
@@
Figure 8: An illustration of the procedure that shows that to each ‘non-planar’ line (black) of a skew-embedded split Cayley hexagon of order two (top) one can associate a unique linear doily (bottom) that shares with it the maximum possible number of lines. The horizontal highlighted line of the hexagon is its axis; the meaning of the remaining colored lines is explained in the text.
@@
The above-described relation can also help us understand why a classically-embedded hexagon and a doily share just three lines belonging to a grid of the doily. For if we disregard in Figure 8 the three colored lines that occur only in skew-embedded hexagons, the remaining three lines (i. e. the black line and the two gray lines) indeed belong to a particular grid of the doily in question!
@@
By a computer search we have found out that there are only three more patterns of lines a linear doily can share with a skew-embedded hexagon. One of them is an already mentioned set of three mutually disjoint lines that belong to some grid of the doily. There exists another three-line pattern, namely that comprising two disjoint lines having a common transversal; slightly rephrased, this pattern entails any three lines forming sides of a quadrangle in the doily. The remaining type features two intersecting lines. These can readily be illustrated employing the copy of skew-embedded hexagon depicted in Figure 9. To this end, let us consider three particular linear three-qubit doilies, namely the doilies whose all 15 observables feature $I$ on the first qubit (the ‘left’ doily), on the second qubit (the ‘middle’ doily) and on the third qubit (the ‘right’ doily). From Figure 9 one can easily discern that the ‘left’ doily (red) shares with the hexagon the following three lines $IXX-IZZ-IYY$, $IYY-IZX-IXZ$ and $IXZ-IIZ-IXI$, the middle one being indeed incident with either of the remaining two that are disjoint. On the other hand, the ‘right’ doily (blue) has two concurrent lines in common with the hexagon: $XZI-IZI-XII$ and $IZI-YZI-YII$. For the sake of completeness, we also mention that the ‘middle’ doily (not shown) shares with the hexagon the maximal pattern, composed of the five lines $ZIY-IIY-ZII$, $YII-IIX-YIX$, $ZIX-XIY-YIZ$, $XIX-YIY-ZIZ$ and $XIZ-IIZ-XII$ forming a spread, and the line $ZII-IIX-ZIX$.
@@
Figure 9: A different copy of skew-embedded hexagon (whose axis is the line $YYZ-IXY-YZX$) employed to illustrate additional doily-hexagon intersection patterns. The points and (hexagon-shared) lines of the two doilies are highlighted in different colors, the three points that belong to both the doilies being colored gray.
@@
The above-described intersection patterns of linear doilies with skew-embedded hexagons are of crucial importance for a better understanding of the fact why the complements of the latter are contextual. This basically boils down to the fact that a less symmetric, skew-embedded hexagon just fits into $\mathcal{W}(5,2)$ in such a manner that it provides, in contrast to its much more symmetric, classically-embedded sibling, enough ‘space’ for its complement to contain grids (aka Mermin-Peres magic squares). To see this explicitly, let us consider a doily having with a skew-embedded hexagon two concurrent lines in common. It is easy to see that out of the ten grids in this doily, there are four of them that do not contain any of the two shared lines. That is, these four grids are fully contained in the complement of the hexagon. A similar situation also occurs in the above-described three-line pattern; in this case the corresponding doily has just two grids devoid any of the three lines and, so, lying fully in the hexagon’s complement. As grids are the simplest, and so fundamental, quantum contextual structures in multi-qubit symplectic polar spaces, their occurrence in the complements of skew-embedded hexagons lends itself as one of the most natural justifications why these complements are contextual.
@@ 4 The contextuality of $\mathcal{W}(5,2)$ and the classical embedding of $\mathcal{H}$
Recently, the authors of [29] made the key (computer-based) discovery that the degree of contextuality of the configuration comprising all 315 line-contexts of $\mathcal{W}(5,2)$ is equal to $63$, and not $90$ (= the number of negative lines) as previously proposed by [7]. In addition, and more importantly, they found out that the 63 unsatisfiable contexts of this configuration are in bijection with 63 lines of a copy of $\mathcal{H}$ that is embedded classically into $\mathcal{W}(5,2)$. These facts came as a big surprise to us and prompted us to have a closer look at what is going on here. We will provide first some algebraic-geometrical arguments for the occurrence of the number $63$. Then, employing a line-layered decomposition of the hexagon, we will demonstrate on several examples how the contextuality properties of a three-qubit configuration having three-element contexts can simply be read off from its generic intersection with a classically-embedded copy of $\mathcal{H}$.
4.1 The degree of contextuality of $\mathcal{W}(5,2)$s sixty-three
@@ 4.2 Layering of the split Cayley hexagon of order two
Figure 10: An illustration of the line-distribution (i. e., the distribution of contexts) within the hexagon.
To see finer traits of the structure of the hexagon, one can redraw the three main ‘layers’ or ‘domains’ of the hexagon, namely red (top left), blue (top right) and yellow (bottom) in the most symmetric way. Note that both the red and blue domains can each be viewed as a pair of circumscribed octagons, either of them being isomorphic to the same $(24_{2},16_{3})$-configuration. In the illustration of the yellow domain each observable is represented by two different (opposite) points and the corresponding affine part of each yellow line, i. e. that two-point part of the line that is left after the removal of the point lying on the axis, has four distinct images forming a quadrangle. Our option for such a rendering of the yellow layer is simple: if one stacks all the three figures above each other then the corresponding three observables at a given position define a particular gray line of the hexagon; for example, the three topmost observables form the gray line $XXI-ZYX-YZX$.
@@ 4.3 A simple recipe how to get from a skew-embedded hexagon a classically-embedded one, and vice versa
The above-described layered structure of the hexagon turns out to be very relevant to better understand the relation between the two types of embeddings of the split Cayley hexagon of order two. In fact, there exists a rather simple recipe that ‘transforms’ a skew hexagon into a classical one. Let us start with the skew hexagon shown in Figure 10. Keep the three observables of the axis intact. On each yellow line, swap the two remaining observables. On each red (blue) line leave each observable intact and on each blue (red) line replace its observable by that which is the product of a swapped yellow observable and a fixed red (blue) observable on the corresponding gray line. What we get is a copy of a classically-embedded hexagon, depicted in Figure 11, that shares 39 lines with the original skew hexagon: the black line, the 6 yellow lines, the 16 blue lines and the 16 red lines.
@@
This construction only works if the black (reference) line is the axis of the skew hexagon. Obviously, we can reverse the process and start with a classically-embedded hexagon to get a skew one. In this case any of its 63 lines can be taken as the reference black line, so we get 63 different skew copies from a given classical one. As there are 120 different classical copies of the hexagon in $\mathcal{W}(5,2)$, the above property also implies that there are as many as $120\times 63=7560$ skew-embedded hexagons living in $\mathcal{W}(5,2)$ – confirming the result of [18] based on an exhaustive computer search.
4.4 Classically-embedded hexagons encode three-qubit contextuality